After deciding to buy a new Mercedes-Benz C Class sedan, you can either lease the car or purchase it on a three-year loan. The car you wish to buy costs $35,000. The dealer has a special leasing arrangement where you pay $99 today and $499 per month for the next three years. If you purchase the car, you will pay it off in monthly payments over the next three years at a 6 percent APR compounded monthly. You believe you will be able to sell the car for $23,000 in three years. Should you buy or lease the car? What break-even resale price in three years would make you indifferent between buying and leasing?
Answer #1


solution: monthlly rate = 0.06/12 number of period = 3*12 = 36 Present value of Purchasing car: Purchase cost Less : Present